# Randomized Algorithms for Low-Rank Matrix and Tensor Decompositions

### v1

@@ 3.3 Application: Randomized DRMs for Overdetermined Least Squares
A natural idea is to apply a randomized DRM to $\bm{\mathsf{A}}$ and $\bm{\mathsf{b}}$ in (28), a strategy known as “sketch-and-solve” sarlos2006improved. Namely, we draw a randomized DRM $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{d\times m}$ and solve a smaller sketched problem
@@
If $\bm{\mathsf{\Gamma}}$ is a randomized subspace embedding for the $k$-dimensional column space of $\begin{bmatrix}\bm{\mathsf{A}}&\bm{\mathsf{b}}\end{bmatrix}$ with distortion $\varepsilon\in(0,1)$, then
@@
Randomized DRMs are also utilized in other methods of solving (28). Sketched GMRES (NakatsukasaTropp2024, Algorithm 1.1) includes a $k$-truncated Arnoldi process to quickly compute an approximate basis of the Krylov subspace. Sketch-and-precondition methods are also commonly used, but we note that these methods were recently shown to be numerically unstable when initialized with the zero vector, cf. MeierNakatsukasaTownsendWebb2024. With iterative sketching methods, sketched least squares problems are solved until converged, and these methods were recently proven to be forward stable in Epperly2024. We discuss alternative randomized DRMs for overdetermined least squares problems for computing tensor decompositions in Sect. 6, which are based on the randomized algorithms for low-rank matrix decompositions that we discuss next.
@@ 4 Randomized Algorithms for Low-Rank Matrix Decompositions
In this section, we illustrate several ways that randomized DRMs are often used to compute low-rank matrix decompositions. We first introduce the rangefinder problem and show how it can be solved with randomized DRMs in Sect. 4.1. We then explain how randomized DRMS can be used to estimate the error in the corresponding low-rank approximation, which can be utilized for adaptive rangefinding. These ideas are foundational for randomized algorithms that have become ubiquitous for fast and reliable low-rank matrix approximation: the randomized SVD (Sect. 4.2) and randomized ID/CUR (Sect. 4.3).
@@ 4.1 The Randomized Rangefinder
At the heart of many linear algebraic tasks lies the rangefinder problem, the computation of a lower-dimensional space that sufficiently captures the action of a given matrix. More specifically, to solve the rangefinder problem for an input matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ and target rank $k\leq\min(m,n)$, we compute an orthonormal matrix $\bm{\mathsf{Q}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ whose columns approximately span the column space of $\bm{\mathsf{A}}$. We then an obtain a rank-$k$ approximation given by $\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$, with approximation error
@@
measured in either the Frobenius or spectral norm in our work. The Frobenius norm is often preferable for ease of analysis, but error bounds in the Frobenius norm can be less informative than bounds in the spectral norm, cf. (martinsson2020, Remark 2.1).
The rangefinder problem can be handled effectively with randomized sketching as shown in Algorithm 1. This algorithm produces an orthonormal matrix $\bm{\mathsf{Q}}$ via QR factorization of an $m\times\ell$ random sketch of $\bm{\mathsf{A}}$ for desired embedding dimension $\ell$. The total cost of Algorithm 1 includes the cost of simulating the $n\times\ell$ random test matrix, $O(mn\ell)$ operations to form the sketch, and $O(mk^{2})$ operations for the QR factorization. Sparsity or availability of fast matrix-vector primitives with $\bm{\mathsf{A}}$ can lend additional performance improvements.
@@
Implementation details
There are many considerations to be made when implementing the randomized rangefinder algorithm in practice. We summarize the important details here and refer to halko2011; martinsson2020 for more thorough treatments.
First, sketching, as opposed to sampling, should be used whenever possible. Gaussian matrices are popular for easy implementation and strong theoretical results, but structured random matrices (e.g., SparseStack) perform almost as well in practice. Second, the numerical accuracy of the rangefinder algorithm can be improved with power iteration Hestenes1952methods; IpsenWills2006; Saad1986GMRES or Krylov methods IpsenMeyer1998; Saad2022; SimonciniSzyld2003, in addition to numerically stable orthogonalization techniques like Householder reflections or rank-revealing QR Chan1987; Ilse1994.
One last important detail concerns the subspace dimension. If the target rank $k$ of the low-rank approximation to $\bm{\mathsf{A}}$ is known a priori, then a randomized DRM of size $\ell=k+p$ is usually sufficient with a small amount of oversampling $p$, such as $p=5$. (We deal with the situation where $k$ is unknown in Sect. 4.1.) For Gaussian test matrices with slight oversampling, (halko2011, Theorem 10.1) guarantees that the expected error of the randomized rangefinder is comparable to the best rank-$k$ approximation error from the truncated SVD of $\bm{\mathsf{A}}$, so long as its trailing singular values are small. This is generally achievable, for instance, by powering $\bm{\mathsf{A}}$ in the sketch $\bm{\mathsf{Y}}=(\bm{\mathsf{A}}\bm{\mathsf{A}}^{*})^{q}\bm{\mathsf{\Omega}}$ some small number of iterations $q$ RokhlinSzlamTygert2009.
The theoretical results for Gaussian random matrices often furnish the performance heuristics for structured random sketching matrices, which behave similarly to Gaussians in practice despite the more pessimistic bounds that appear in their analysis (halko2011, Theorem 11.2). This claim is strengthened by the simple addition of a posteriori “certificates of accuracy” upon termination of Algorithm 1, as explained in the next section.
@@
We now demonstrate how the approximation error (30) can be estimated reliably and efficiently with randomized sketching. This method of error estimation also serves as a “certificate of accuracy” for rangefinding, an assurance that the orthornomal basis given by $\bm{\mathsf{Q}}$ produces a satisfactory low-rank approximation of $\bm{\mathsf{A}}$.
Suppose that we have an orthonormal matrix $\bm{\mathsf{Q}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ whose columns potentially form a basis for the range of $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, e.g., the output of Algorithm 1 for a given $k$. We estimate the approximation error $\|(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\|$ using the following method.
Let $\bm{\mathsf{\Phi}}\in\mathbb{F}^{\hskip 0.56905ptn\times s}$ be a Gaussian matrix that is independent from any random matrix used to compute $\bm{\mathsf{Q}}$. In practice, $s$ is a small fixed value, such as $s=10$. We form a new random sketch, known as the auxiliary sample
@@
incurring only $s$ additional matrix-vector products with $\bm{\mathsf{A}}$. Using $\bm{\mathsf{Z}}$, we can obtain an inexpensive Gaussian sketch of the approximation error (30):
@@
We then take advantage of randomized methods to estimate the norm of (32), such as randomized trace estimation for the Frobenius norm or randomized Krylov methods for the Schatten $2p$-norm (a proxy for the the spectral norm with sufficiently large $p$). Recall that the Schatten $p$-norm of a matrix is the $\ell_{p}$-norm of its singular values, so that the Schatten $\infty$-norm coincides with the spectral norm and the Schatten 2-norm with the Frobenius norm.
To estimate the error in the Frobenius norm, we use the approximation
@@
relying on the theory of randomized trace estimation (recall that $\textup{trace}(\bm{\mathsf{X}}^{*}\bm{\mathsf{X}})=\|\bm{\mathsf{X}}\|_{\text{F}}^{2}$). Namely, we exploit the fact that for any matrix $\bm{\mathsf{B}}$ and isotropic random vector $\omega$, $\text{trace}(\bm{\mathsf{B}})=\bm{\mathsf{\omega}}^{*}\bm{\mathsf{B}}\bm{\mathsf{\omega}}$ in expectation. To reduce the variance, we take the average of $s$ random variables in (33), using the auxiliary sample $\bm{\mathsf{Z}}$.
Estimating the error in the spectral norm is not quite as simple. It is possible to construct unbiased estimators for the Schatten 2pp-norm for any positive integer $p$, as for the Frobenius norm above, but these estimators are expensive with large variance KongValiant2017; LiWoodruff2014. Randomized Krylov methods are often used instead to estimate the spectral norm LeykWoz1998; MartinssonRokhlinTygert2011; MuscoMusco2015; Tropp2022; WangZhangZhang2015. For example, we can estimate the spectral norm error with a randomized block Krylov method, letting $\bm{\mathsf{B}}=\begin{bmatrix}(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\end{bmatrix}\begin{bmatrix}(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\end{bmatrix}^{*}$,
$\|(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\|^{2}_{2}\approx\max_{\bm{\mathsf{v}}\in\mathcal{K}(\bm{\mathsf{B}};\bm{\mathsf{\Phi}})}\frac{\bm{\mathsf{v}}^{*}\bm{\mathsf{B}}\bm{\mathsf{v}}}{\bm{\mathsf{v}}^{*}\bm{\mathsf{v}}},$ (34)
where $\mathcal{K}(\bm{\mathsf{B}};\bm{\mathsf{\Phi}})=\textup{span}\begin{bmatrix}\bm{\mathsf{\Phi}}\ &\bm{\mathsf{B}}\bm{\mathsf{\Phi}}\ &\bm{\mathsf{B}}^{2}\bm{\mathsf{\Phi}}\ &\cdots&\bm{\mathsf{B}}^{q}\bm{\mathsf{\Phi}}\end{bmatrix}$ is a block Krylov subspace of depth $q$. In general, $q\approx\log m$ is necessary to achieve the fast convergence rate of $1/q^{2}$ for Krylov methods, but smaller values of $q$ can be used in practice for matrices with fast spectral decay. Although Krylov methods enjoy quick convergence, numerical instability is a known shortcoming, so careful implementation is needed, cf. MuscoMusco2015; Tropp2018; WangZhangZhang2015.
Regardless of the choice of norm, an auxiliary sketch of the error with a Gaussian random test matrix $\bm{\mathsf{\Phi}}$ can provide us with an a posteriori “certificate of accuracy” for rangefinding (martinsson2020, Section 12.2). For instance, if a structured random matrix (e.g., SRTT, sparse sign) is used in Algorithm 1 instead of a Gaussian matrix, the computations are generally faster but lack formal guarantees on the accuracy of the computed solution. More generally, if we are unsure of the accuracy of a given solution to the rangefinder problem, we can take a small (independent) Gaussian sketch of the error as in (31). This auxiliary sketch is inexpensive and avails us of the powerful theory of Gaussian matrices, regardless of how the column space approximation was obtained. In the next section, we demonstrate how this method of error estimation can be utilized for adaptive rangefinding when the target rank is unknown a priori.
Adaptive rangefinding
Up to this point, we have assumed that the target approximation rank could be passed as an argument to the randomized rangefinder algorithm. Unfortunately, a satisfactory target rank for the approximation is frequently unknown a priori. In that case, we can implement the randomized rangefinder algorithm adaptively, using the error estimation techniques of the previous section to approximate the range of $\bm{\mathsf{A}}$ to a desired precision $\varepsilon$, forming $\bm{\mathsf{Q}}$ incrementally until that precision is achieved.
For ease of notation, let norm_est($\bm{\mathsf{Z}}$) denote the rangefinder error estimate obtained by one of the techniques in Sect. 4.1, e.g., (33). An adaptive version of the randomized rangefinder is given in Algorithm 2, which we quickly summarize.
@@
7:while $\texttt{norm\_est}(\bm{\mathsf{Y}})>\varepsilon$ do $\triangleright$ Error estimation of Sect. 4.1
@@
Let $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ and let $b$ be the number of columns to be processed at a time. The choice of block size $b$ affects the computational efficiency of the algorithm. If $b$ is too small, we are not fully reaping the benefits of blocked matrix operations that help reduce memory traffic and avoid communication bottlenecks Heavner2022; Orti1998. With a larger block size $b$, more floating-point operations can be cast as matrix-matrix multiplication (BLAS-3) vs. matrix-vector multiplication (BLAS-2) Blackford2002; Dongarra1990; Gallivan1987; Higham1990. However, a larger block size runs the risk of overshooting the target rank, meaning we have done more work than necessary for rangefinding. In practice, a block size $10\leq b\leq 100$ is usually appropriate.
In each iteration of the while-loop in Algorithm 2, the matrix $\bm{\mathsf{Y}}$ holds a sample of the residual error after the execution of line 9:
@@
where $\bm{\mathsf{Q}}$ is the basis that has been built up to the current iteration. Critically, the random test matrix $\bm{\mathsf{\Omega}}$ used in (35) is independent from the random matrices used to compute the current basis $\bm{\mathsf{Q}}$. We note that to compute the very first residual error norm_est($\bm{\mathsf{Y}}$) in line 5, we can draw a small auxiliary sample that then can be recycled in the second batch of samples in line 8.
Variations of Algorithm 2 have been designed to optimize performance for specific types of input matrices. If the input matrix $\bm{\mathsf{A}}$ is small enough to store in RAM, it can be explicitly updated in each iteration via $\bm{\mathsf{A}}=(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}$ as in (martinsson2020, Algorithm 14). Alternatively, the version of YuGuLiLiuLi2017 effectively handles input matrices stored out-of-core, as well as sparse input matrices. If the spectrum of the input matrix decays slowly, power iteration can also be incorporated into Algorithm 2 to improve performance MartinssonVoronin2016.
@@ 4.2 Randomized SVD
Now equipped with the randomized rangefinder, we present a foundational algorithm for low-rank matrix approximation known as the randomized SVD halko2011. Pseudocode is provided in Algorithm 3. We quickly go through its steps to show that the error of the resulting approximate SVD is precisely the randomized rangefinder error.
@@
Given an input matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, our goal is to compute an approximate rank-$k$ truncated singular value decomposition
$\bm{\mathsf{A}}\approx\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*},$ (36)
where $\bm{\mathsf{U}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ and $\bm{\mathsf{V}}\in\mathbb{F}^{\hskip 0.56905ptn\times k}$ are orthonormal matrices, and $\bm{\mathsf{\Sigma}}\in\mathbb{F}^{\hskip 0.56905ptk\times k}$ is a diagonal matrix of entries approximating the largest singular values of $\bm{\mathsf{A}}$.
Suppose that $\bm{\mathsf{Q}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ is a solution to the rangefinder problem (not necessarily from Algorithm 1), which corresponds to a rank-$k$ approximation of $\bm{\mathsf{A}}$ given by $\bm{\mathsf{A}}\approx\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$. Let $\bm{\mathsf{B}}=\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptk\times n}$ and compute singular value decomposition
$\bm{\mathsf{B}}=\underset{k\times k}{\widehat{\bm{\mathsf{U}}}}\ \underset{k\times k}{\bm{\mathsf{\Sigma}}}\ \underset{k\times n}{\bm{\mathsf{V}}^{*}}.$ (37)
We observe that
$\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}\|$ $=\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\bm{\mathsf{B}}\|=\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\widehat{\bm{\mathsf{U}}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}\|=\|\bm{\mathsf{A}}-\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}\|,$ (38)
for $\bm{\mathsf{U}}:=\bm{\mathsf{Q}}\widehat{\bm{\mathsf{U}}}$; i.e. the approximation error of $\bm{\mathsf{A}}\approx\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}$ is exactly the approximation error of $\bm{\mathsf{A}}\approx\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$.
Thus, if $\bm{\mathsf{Q}}=\texttt{RandomizedRangefinder}(\bm{\mathsf{A}},k)$ is the rangefinder solution computed by Algorithm 1, the accuracy of the randomized SVD is equal to that of the randomized rangefinder. In particular, this means that the error estimation techniques of Sect. 4.1 can also be used to estimate the error of the randomized SVD. Then if the target rank $k$ is unknown a priori, the adaptive randomized rangefinder (Algorithm 2) can be carried out first, which determines $k$ in the randomized SVD.
The total cost of Algorithm 3 includes the cost of the randomized rangefinder, plus $O(mnk)$ floating-point operations for $\bm{\mathsf{B}}=\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$ in line 2. We note that the cost of the SVD in line 3 is just $O(k^{2}n)$, so the cost of Algorithm 3 after rangefinding is usually dominated by the matrix-matrix multiplication. Executing Algorithm 3 requires storage for $O((m+n)k)$ numbers.
Recent work has also been done on cost-effective a posteriori error estimation methods that use only the information at hand for computing the low-rank decomposition. In EpperlyTropp2024, an algorithm for a posteriori error estimation is presented for the randomized SVD in a matrix-free setting, where the input matrix $\bm{\mathsf{A}}$ is only accessible via mat-vecs. Given the rank-$k$ randomized SVD $\bm{\mathsf{X}}=\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}\in\mathbb{F}^{m\times n}$ corresponding to Gaussian random matrix $\bm{\mathsf{\Omega}}=\begin{bmatrix}\bm{\mathsf{\omega}}_{1}&\bm{\mathsf{\omega}}_{2}&\ldots&\bm{\mathsf{\omega}}_{k}\end{bmatrix}$, we want to estimate the mean-square error
@@
where $\bm{\mathsf{X}}_{k-1}\in\mathbb{F}^{m\times(k-1)}$ is a rank-(k1)(k-1) randomized SVD.
By (EpperlyTropp2024, Theorem 2.1), the following is an unbiased estimator for $\textup{MSE}(\bm{\mathsf{X}}_{k-1},\bm{\mathsf{A}})$, which we use as a proxy for the mean-square error of the rank-$k$ approximation $\bm{\mathsf{X}}$, called the leave-one-out error estimator:
@@
where $\bm{\mathsf{X}}^{(j)}=\bm{\mathsf{Q}}^{(j)}(\bm{\mathsf{Q}}^{(j)})^{*}\bm{\mathsf{A}}$ is called a replicate, formed from the matrix $\bm{\mathsf{Q}}^{(j)}$ outputted by the randomized rangefinder using $\bm{\mathsf{\Omega}}$ without its $j$-th column $\bm{\mathsf{\omega}}_{j}$. Implemented naively, the right-hand side of (39) is too expensive for practical use, as it would require $k$ additional randomized rangefinder applications. However, by (EpperlyTropp2024, Equation 4.2),
@@
where $\bm{\mathsf{t}}_{j}$, $j=1,\ldots,k$, are the normalized columns of $(\bm{\mathsf{R}}^{*})^{-1}$ in the QR factorization defined by $\bm{\mathsf{Q}}$ in step 1 of the randomized SVD (now also storing the $\bm{\mathsf{R}}$ matrix). Then we can obtain each replicate $\bm{\mathsf{X}}^{(j)}$ inexpensively as
$\bm{\mathsf{X}}^{(j)}=\bm{\mathsf{U}}(\bm{\mathsf{I}}-\widehat{\bm{\mathsf{U}}}^{*}\bm{\mathsf{t}}_{j}\bm{\mathsf{t}}_{j}^{*}\widehat{\bm{\mathsf{U}}})\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*},$ (41)
which allows us to compute (39) without any additional mat-vecs or applications of Algorithm 3, via

### v2

@@ 3.3 Application: Randomized DRMs for Overdetermined Least Squares
A natural idea is to apply a randomized DRM to $\bm{\mathsf{A}}$ and $\bm{\mathsf{b}}$ in (27), a strategy known as “sketch-and-solve” sarlos2006improved. Namely, we draw a randomized DRM $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{d\times m}$ and solve
@@
If $\bm{\mathsf{\Gamma}}$ embeds the $k$-dimensional column space of $\begin{bmatrix}\bm{\mathsf{A}}&\bm{\mathsf{b}}\end{bmatrix}$ with distortion $\varepsilon\in(0,1)$, then
@@
Randomized DRMs are also utilized in other methods of solving (27). Sketched GMRES (NakatsukasaTropp2024, Algorithm 1.1) includes a $k$-truncated Arnoldi process to quickly compute an approximate basis of the Krylov subspace. Sketch-and-precondition methods are also popular Avron2010; MeierNakatsukasaTownsendWebb2024; MengSaundersMahoney2014. We discuss alternative randomized DRMs for overdetermined least squares problems in computing tensor decompositions in Sect. 6.
@@ 4 Randomized Algorithms for Low-Rank Matrix Decompositions
In this section, we illustrate several ways that randomized DRMs are often used to compute low-rank matrix decompositions. We introduce the rangefinder problem and show how it can be solved with randomized DRMs in Sect. 4.1. We then explain how randomized DRMS can be used to estimate the error in the corresponding low-rank approximation for adaptive rangefinding. These ideas are foundational for randomized algorithms that have become ubiquitous for fast and reliable low-rank matrix approximation: the randomized SVD (Sect. 4.2) and randomized ID/CUR (Sect. 4.3).
@@ 4.1 The Randomized Rangefinder
At the heart of many linear algebraic tasks lies the rangefinder problem, finding a lower-dimensional space that captures the action of a given matrix. To do this, for an input matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ and target rank $k\leq\min(m,n)$, we compute an orthonormal matrix $\bm{\mathsf{Q}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ whose columns approximately span the column space of $\bm{\mathsf{A}}$. We then obtain a rank-$k$ approximation given by $\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$, with approximation error
@@
measured in either the Frobenius or spectral norm in our work. The Frobenius norm is often preferable for ease of analysis but can be less informative the spectral norm.
The rangefinder problem can be handled effectively with randomized DRMs, as in Algorithm 1, which produces an orthonormal matrix $\bm{\mathsf{Q}}$ via QR factorization of an $m\times\ell$ random sketch of $\bm{\mathsf{A}}$ for desired embedding dimension $\ell$. The total cost of Algorithm 1 includes the cost of simulating the $n\times\ell$ random test matrix, $O(mn\ell)$ operations to form the sketch, and $O(mk^{2})$ operations for the QR factorization. Sparsity or availability of fast matrix-vector primitives with $\bm{\mathsf{A}}$ can lend additional performance improvements.
@@
We mention a few implementation details here and refer to halko2011; martinsson2020 for thorough treatments. For Gaussian sketching matrices with dimension $k+p$, (halko2011, Theorem 10.1) guarantees the expected error of the randomized rangefinder is comparable to the best rank-$k$ approximation error from the truncated SVD, for very small $p$, e.g., $p=5$, so long as the trailing singular values of $\bm{\mathsf{A}}$ are small. The theory for Gaussians often furnishes performance heuristics for other randomized DRMs, which frequently behave like Gaussians despite the more pessimistic theoretical bounds, cf. Table 1. The simple addition of a posteriori “certificates of accuracy” provides additional assurance.
@@
The approximation error (29) can be estimated reliably and efficiently with randomized sketching, which serves as a “certificate of accuracy” that the basis given by $\bm{\mathsf{Q}}$ approximately spans the range of $\bm{\mathsf{A}}$. Let $\bm{\mathsf{Q}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ be an orthonormal matrix whose columns potentially form a basis for the range of $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$. Let $\bm{\mathsf{\Phi}}\in\mathbb{F}^{\hskip 0.56905ptn\times s}$ be a Gaussian matrix that is independent from any random matrix used to compute $\bm{\mathsf{Q}}$. In practice, $s$ is a small fixed value, such as $s=10$. We form a sketch, often called the auxiliary sample,
@@
From $\bm{\mathsf{Z}}$, we obtain an inexpensive Gaussian sketch of the approximation error (29):
@@
We can then estimate the Frobenius norm of (31) as
@@
noting that $\text{trace}(\bm{\mathsf{B}})=\bm{\mathsf{\omega}}^{*}\bm{\mathsf{B}}\bm{\mathsf{\omega}}$ in expectation for any matrix $\bm{\mathsf{B}}$ and isotropic random vector $\omega$). Then to reduce the variance, we take the average of $s$ random variables in (32), using the auxiliary sample $\bm{\mathsf{Z}}$. The error sketch avails us of the powerful theory of Gaussian matrices, regardless of how the column space approximation was obtained. For spectral norm estimation, see (martinsson2020, Sects. 5-6).
Rank-adaptive rangefinding
We have so far assumed the target approximation rank can be passed as an argument to the randomized rangefinder. Unfortunately, the target rank for an approximation is often unknown a priori. The method of error estimation above also enables a rank-adaptive randomized rangefinder algorithm (Algorithm 2). For ease of notation, let norm_est($\bm{\mathsf{Z}}$) denote the rangefinder error estimated as in (32).
@@
7:while $\texttt{norm\_est}(\bm{\mathsf{Y}})>\varepsilon$ do $\triangleright$ Error estimation as in (32)
@@
Let $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ and let $b$ be the number of columns processed at a time. In practice, $10\leq b\leq 100$ is usually appropriate. In each iteration of Algorithm 2, after line 9,
@@
where $\bm{\mathsf{Q}}$ is the currently constructed basis. Critically, the random test matrix $\bm{\mathsf{\Omega}}$ in (33) is independent of the random matrices used to compute $\bm{\mathsf{Q}}$.
@@ 4.2 Randomized SVD
The randomized SVD of halko2011 (Algorithm 3) is an important application of rangefinding, with approximation error equal to the randomized rangefinder error.
@@
Given an input matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, our goal is to compute an approximate rank-$k$ truncated singular value decomposition $\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}.$ Suppose that $\bm{\mathsf{Q}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ is the output of Algorithm 1, corresponding to rank-$k$ approximation $\bm{\mathsf{A}}\approx\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$. Let $\bm{\mathsf{B}}=\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptk\times n}$ and compute its SVD $\bm{\mathsf{B}}=\widehat{\bm{\mathsf{U}}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}$. We observe that for $\bm{\mathsf{U}}:=\bm{\mathsf{Q}}\widehat{\bm{\mathsf{U}}}$,
$\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}\|$ $=\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\bm{\mathsf{B}}\|=\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\widehat{\bm{\mathsf{U}}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}\|=\|\bm{\mathsf{A}}-\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}\|.$ (34)
Thus, the approximation error of $\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}$ is exactly the approximation error of $\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$. If the target rank $k$ is unknown a priori, Algorithm 2 can be used instead.
The total cost of Algorithm 3 includes the cost of rangefinding, plus $O(mnk)$ floating-point operations for $\bm{\mathsf{B}}=\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$ in line 2. We note the cost of the SVD in line 3 is only $O(k^{2}n)$. The storage cost is $O((m+n)k)$.
The recent work of EpperlyTropp2024 introduces a cost-effective a posteriori error estimation method for the randomized SVD. Given rank-$k$ randomized SVD $\bm{\mathsf{X}}=\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}\in\mathbb{F}^{m\times n}$ corresponding to Gaussian DRM $\bm{\mathsf{\Omega}}=\begin{bmatrix}\bm{\mathsf{\omega}}_{1}&\ldots&\bm{\mathsf{\omega}}_{k}\end{bmatrix}$, the mean-square error is defined
@@
where $\bm{\mathsf{X}}_{k-1}$ is a rank-(k1)(k-1) randomized SVD. By (EpperlyTropp2024, Theorem 2.1), the following is an unbiased estimator for $\textup{MSE}(\bm{\mathsf{X}}_{k-1},\bm{\mathsf{A}})$, called the leave-one-out error estimator:
@@
where $\bm{\mathsf{X}}^{(j)}=\bm{\mathsf{Q}}^{(j)}(\bm{\mathsf{Q}}^{(j)})^{*}\bm{\mathsf{A}}$ is called a replicate, formed from the matrix $\bm{\mathsf{Q}}^{(j)}$ outputted by Algorithm 1 using $\bm{\mathsf{\Omega}}$ without its $j$-th column $\bm{\mathsf{\omega}}_{j}$. Implemented as written, the right-hand side of (35) is too expensive for practical use. However, by (EpperlyTropp2024, Equation 4.2),
@@
where $\bm{\mathsf{t}}_{j}$, $j=1,\ldots,k$, are the normalized columns of $(\bm{\mathsf{R}}^{*})^{-1}$ in the QR factorization defined by $\bm{\mathsf{Q}}$ in step 1 of the randomized SVD (now also storing the $\bm{\mathsf{R}}$ matrix). Then we can obtain each replicate from (36) inexpensively as $\bm{\mathsf{X}}^{(j)}=\bm{\mathsf{U}}(\bm{\mathsf{I}}-\widehat{\bm{\mathsf{U}}}^{*}\bm{\mathsf{t}}_{j}\bm{\mathsf{t}}_{j}^{*}\widehat{\bm{\mathsf{U}}})\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*},$ which allows us to compute (35) without any additional rangefinding via